two integers, a and b, between 1 and 20,such that a squared plus b squared is a perfect square
Answer by Edwin McCravy(20086)




 a2 +  b2 =  c2
------------------------
 32 +  42 =  52
 52 + 122 = 132
 62 +  82 = 102
 82 + 152 = 172
 92 + 122 = 152
122 + 162 = 202
152 + 202 = 252

Edwin